OpenStudy (anonymous):

Find the x- and y-intercepts and graph the equation plotting the intercepts. Please show all of your work. -3x-4y=18

OpenStudy (anonymous):

so its -1

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

?????

OpenStudy (lgbasallote):

x intercept...set y = 0 solve for x y int...set x= 0 solve y graph ok?

OpenStudy (anonymous):

confused

OpenStudy (anonymous):

i dont understand

OpenStudy (lgbasallote):

x intercept...set y to zero so... -3x - 4(0) = 18 -3x - 0 = 18 -3x = 18 --- ----- -3 -3 x = -6 (-6,0) is one point understand now?

OpenStudy (anonymous):

sry i really dont

OpenStudy (anonymous):

-3(0)-4y=18 0-4y=18 -4y=18......but wht do i do from here

OpenStudy (lgbasallote):

divide both sides by -4

OpenStudy (anonymous):

-4 doesnt go into 18

OpenStudy (anonymous):

-4.5?

OpenStudy (lgbasallote):

yes...so it's (0.-4.5) now graph